Nail polish remover is essentially a mixture of water and a chemical called acetone. How much pure acetone must be combined with a solution that is 19% acetone to make 27 oz of a 37% solution
please show me how to do this

Found 2 solutions by Alan3354, macston:
Answer by Alan3354(69443)

Nail polish remover is essentially a mixture of water and a chemical called acetone. How much pure acetone must be combined with a solution that is 19% acetone to make 27 oz of a 37% solution
======================
Add A ounces to the 19% solution (S)
---
A + S = 27 (total solution)
100*A + 19*S = 27*37 (total acetone)
S = 27 - A
---
100A + 19(27-A) = 999
----
Solve for A

Answer by macston(5194)

To have 37% acetone of 27 oz. leaves 63% water. 0.63*27 oz=17.01 oz water
A 19% solution is 81% water so 17.01/.81=21 oz
We need 21 oz 0f 19% solution for the water to make 27 oz of 37% solution.
To 21 oz of 19% solution, we add 6 oz acetone to make 27 oz. of 37% solution.